A sample of naturally occurring silicon consists Si-28 (amu=27.9769),Si-29 (amu=28.9765) and Si-30 (amu=29.9738).If the atomic mass of silicon is 28.0855 and the natural abundance of Si-29 is 4.7%,determine the natural abundances of Si-28 and Si-30
1
Expert's answer
2020-07-24T14:52:13-0400

28.9765 x 0.047 = 1,3618955

28.0855-1,3618955 = 26,7236045

27.9769x + 29.9738y = 26,7236045

x+y = 0,953

y= 0,0186057533 (1.8%)

x= 0,934394247 (93.43%)